slipknot3066:

12(2x - 8) = 4(5 - 2x) + 3(7 - 3x)

dude:

Distribute first, do you need help with that?

ZoeyBiocth16:

expand 12(2x-8): 24x-96 apply distributive law: a(b-c) = ab-ac a=12 b=2x c= 8 12*2x-12*8 simplify: 12*2x-12*8: 24x-96 12*8= 96 multiply 12*2=24: 24x-12*8 24x-96+96= -17x+41+96 simplify: 24x=-17x+137 add 17x to both sides 24x+17x=-17x+137+17x simplify 41x=137 41x/41 = 137/41 x=137/41

slipknot3066:

......Jesus H Christ

ZoeyBiocth16:

lmao

ZoeyBiocth16:

do you get it tho

slipknot3066:

Yea